Tom, Dick & Harry each have a lawn mower and decide to mow a 5900 square foot lawn.
Tom mows 70 square feet per minute, Dick 50 and Harry 40.
Dick and Harry start mowing the lawn at the same time.
Something is wrong with Toms mower and is delayed for 30 mins.
All 3 of them stop mowing at the same time.
How long does Tom mow?
I know the answer is 20, but the route I took to get the answer was a long way. Is there anyway to work this out but in a short way?

Answer by richwmiller(17219)

time for dick and harry t=50
tom=m=t-30
40t+50t+70(t-30)=5900
90t+70t-2100=5900
160t=8000
20t=1000
2t=100
t=50
tom m=50-30=20